Head-First Company now sells both bicycle helmets and motorcycle helmets. Next year, Head- First expects to produce total revenue of $575,000 and incur total variable cost of $400,000. Total fixed cost is expected to be $60,500.

Required:
1. Calculate the break-even point in sales dollars for Head-First. Round the contribution margin ratio to four decimal places and sales to the nearest dollar.
2. Check your answer by preparing a contribution margin income statement.
Answer #1

1)

break even point = fixed cost / contribution margin ratio

contribution = selling price - variable cost

selling price = 575,000

variable cost = 400,000

575,000 - 400,000 = 175,000

contribution margin ratio =( contribution / revenue ) *100

(175,000 / 575,000)*100 = 30.4347 %

break even in dollars = 60,500 / 30.4347%

break even in dollar = 198,786

2) contribution income statement

sales 575,000
product cost
variable cost 400,000
contribution margin 175,000
period cost
fixed cost 60,500
operating income 114,500